OpenStudy (anonymous):

Alex has been saving quarters and dimes. He found that he had 48 coins worth $9.50. How many dimes and quarters did he have?

OpenStudy (jjuden):

is there multiple answers?

OpenStudy (anonymous):

no

OpenStudy (bibby):

represent the information using 2 linear equations

OpenStudy (anonymous):

what 2 linear equations?

OpenStudy (bibby):

for example, if we had 30 coins x+y=30

OpenStudy (bibby):

and then another equation for the sum of the coins

OpenStudy (anonymous):

ok, I understand how to write the first equations, but I'm not sure about the second one for the sum of coins

OpenStudy (bibby):

you need to decide which variable presents which coin then multiply each coin by the value (0.10, 0.25, ,etc.)

OpenStudy (anonymous):

I got .25x+.10y=9.50 for the second equation. is that right?

OpenStudy (bibby):

that looks good, nice job this means that x is the quarters and y is the dimes

OpenStudy (anonymous):

then how do I solve the system of equations?

OpenStudy (bibby):

I don't remember any methods besides substitution x+y=48 .25x+.10y=9.50 y=48-x \(.25x+(.10*(48-x))=9.50\\ .25x+(4.8-0.1x)=9.50\\\)

OpenStudy (anonymous):

thank you!

OpenStudy (bibby):

I think I want to try it the other way around because I get a repeating decimal this way. x+y=48 .25x+.10y=9.50 x=48-y .25(48-y)+.10y=9.50 12-.25y+.10y=9.50 12-.15y=9.50 -.15y=-2.50 .15y=2.50 y=16.6667 looks like you can't avoid it

OpenStudy (bibby):

I'm pretty sure I messed something up here

OpenStudy (bibby):

but somehow we're supposed to have 2/3rds of a dime http://www.wolframalpha.com/input/?i=x%2By%3D48%3B%20.25x%2B.10y%3D9.50%20&t=crmtb01